# AutoRev: Multi-Modal Graph Retrieval for Automated Peer-Review Generation

### v2

@@
Abstract.
Enhancing the quality and efficiency of academic publishing is critical for both authors and reviewers, as research papers are central to scholarly communication and a major source of high-quality content on the web. To support this goal, we propose AutoRev, an automatic peer-review system designed to provide actionable, high-quality feedback to both reviewers and authors. AutoRev leverages a novel Multi-Modal Retrieval-Augmented Generation (RAG) framework that combines textual and graphical representations of academic papers. By modelling documents as graphs, AutoRev effectively retrieves the most pertinent information, significantly reducing the input context length for LLMs and thereby enhancing their review generation capabilities. Experimental results show that AutoRev outperforms state-of-the-art baselines by up to 58.72% and demonstrates competitive performance in human evaluations against ground truth reviews. We envision AutoRev as a powerful tool to streamline the peer-review workflow, alleviating challenges and enabling scalable, high-quality scholarly publishing. By guiding both authors and reviewers, AutoRev has the potential to accelerate the dissemination of quality research on the web at a larger scale. Code will be released upon acceptance.
Graph Neural Networks (GNNs), Retrieval Augmented Generation (RAG), Large Language Models (LLMs), Automatic Peer Reviewing
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generationccs: Information systems Language modelsccs: Information systems Document structure
1. Introduction
@@
Figure 1. Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.:
Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers (Yu et al., 2024; Aitymbetov and Zorbas, 2025; D’Arcy et al., 2024; Idahl and Ahmadi, 2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation (Ye et al., 2024). Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
@@
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community (Chitale et al., 2025; Gao et al., 2025; Agarwal et al., 2025; Zhu et al., 2025). Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, in generating reviews that closely replicate the structure and quality of those produced by existing peer review systems in conferences. Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM. Reducing the token length of the input is shown to enhance LLM performance (Hsieh et al., 2024). To summarize our main contributions:
  1. (1)
    We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers. To our knowledge, this is the first approach which uses graphs in this domain.
  2. (2)
    Our approach improves over SoTA baselines by 58.72% on average across all metrics while being computationally efficient.
  3. (3)
    We conduct thorough human evaluations and other automatic evaluations to demonstrate the competitive performance of our method against ground truth reviews.
It is important to highlight that this paper does not seek to determine the final reviews of a paper. Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
We envision AutoRev as a step towards scalable, automated peer-review systems that can complement human judgment. With the continued surge in conference submissions across disciplines, the demand for such tools is increasingly pressing [^1] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/. By alleviating the bottlenecks of the traditional peer-review process, AutoRev can help accelerate the dissemination of high-quality research on the web, ensuring that credible scholarly knowledge remains accessible at scale.
2. Related Works
A significant amount of research has been conducted on aiding peer reviews for academic papers. Some studies explore the involvement of AI in the reviewer selection process (Aitymbetov and Zorbas, 2025), while others focus on generating meta-reviews (Kumar et al., 2023). Prior work has also examined how AI/NLP techniques can contribute to various aspects of the peer-review pipeline (Kuznetsov et al., 2024).
Faizullah et al. (2024) explore a DPR-based approach for generating suggestive limitations, highlighting the necessity of retrieval methods to overcome length constraints. Building on this, we use DPR outputs as the training data for our Graph Neural Network (GNN), enabling it to learn to identify salient passages for review generation. Yu et al. (2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4222https://arxiv.org/abs/2303.08774 consolidation and proposes a self-correction strategy during fine-tuning. Ye et al. (2024) further caution that LLM-generated reviews may overrate papers or produce superficial feedback, reinforcing our design goal of providing a structured and constructive review rather than final accept/reject decisions.
Among the most relevant recent systems, MARG (D’Arcy et al., 2024) employs a multi-agent framework, distributing different sections of a paper to specialized GPT-4 agents. This enables the model to handle long papers and produce more specific comments, but it comes at the cost of heavy computation and frequent inter-agent coordination errors. OpenReviewer (Idahl and Ahmadi, 2025) instead fully fine-tunes Llama-3.1-8B-Instruct on 79k reviews from ICLR and NeurIPS to generate structured reviews, adhering to the conference guidelines. This method of large-scale full fine-tuning is very costly.
We adopt SEA (Yu et al., 2024) as our primary baseline since it is closest in scope to our work, representing a recent state-of-the-art framework for automated peer review generation on the same ICLR dataset. Other systems such as MARG and OpenReviewer differ substantially in design. MARG uses multi-agent GPT-4 orchestration, while OpenReviewer relies on large-scale fine-tuning, making direct benchmarking less meaningful. Thus, we benchmark against SEA for a fair and reproducible comparison.
@@ 2. Related Works
The integration of graph-based techniques into Retrieval Augmented Generation (RAG) (Lewis et al., 2020) is an emerging research area that aims to enhance the reasoning and factual grounding of LLMs. A recent comprehensive survey by Zhu et al. (2025) provides a foundational taxonomy for this field, categorizing systems based on their database construction, retrieval algorithms, and processing pipelines. Within this framework, AutoRev is a domain-specific Graph-RAG application that generates a hierarchical graph directly from a scientific paper’s text. It employs a learning-based Graph Attention Network (GAT) (Veličković et al., 2018) for retrieval that reduces the context length of input to LLMs, reducing the computational costs.
3. Dataset
@@
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4–consolidated reviews for NeurIPS 2023 are not released. Hence, our analysis focuses only on the ICLR 2024 papers. As reported in Yu et al. (2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
Table 1. Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
@@
Each paper has multiple peer reviews including sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating to assess the quality and clarity of the paper.
From the dataset, we utilize only the Summary, Strengths, Weaknesses, and Questions sections from each review, along with the full content of the corresponding paper. Other sections, such as Ratings are excluded, as our framework is designed to provide constructive feedback rather than produce a final decision. The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers. For our finetuning process, we utilize the GPT-4 consolidated reviews dataset introduced by Yu et al. (2024). This dataset is constructed by using GPT-4 to consolidate multiple human-written peer reviews from OpenReview [^3] https://openreview.net/ into a single, unified review for each paper.
4. Methodology
Figure 2. The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to review generation. The trained GNN selects important passages without requiring human-generated reviews. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final review efficiently, without external grounding.:
In this section, we describe the implementation of our proposed framework. Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) (Karpukhin et al., 2020). The extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews, as discussed in Section 4.4. Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5. An overview of our methodology can be found in Figure 2.
4.1. DPR-based Passage Extraction
Dense Passage Retrieval (DPR) is a neural retrieval model, primarily trained on question-answering datasets, that encodes both queries and context into dense vector representations. Given an input query, DPR retrieves the most relevant passages based on their semantic similarity in the dense embedding space. We use a pre-trained DPR model, to identify the most relevant passages corresponding to the different sections of the reviews that are Summary, Strengths, Weaknesses, and Questions, for each paper in the training set. Specifically, we use the context encoder [^4] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and the query encoder [^5] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.

### v3

# Graph-Guided Passage Retrieval for Author-Centric Structured Feedback

@@
Abstract
Obtaining high-quality, pre-submission feedback is a critical bottleneck in the academic publication lifecycle for researchers. We introduce AutoRev, an automated author-centric feedback system that generates structured, actionable guidance prior to formal peer review. AutoRev employs a graph-based retrieval-augmented generation framework that models each paper as a hierarchical document graph, integrating textual and structural representations to retrieve salient content efficiently. By leveraging graph-based passage retrieval, AutoRev substantially reduces LLM input context length, leading to higher-quality feedback generation. Experimental results demonstrate that AutoRev significantly outperforms baselines across multiple automatic evaluation metrics, while achieving strong performance in human evaluations. Code will be released upon acceptance.
Graph-Guided Passage Retrieval for Author-Centric Structured Feedback
Maitreya Prafulla Chitale1thanks: Equal contribution.   Ketaki Mangesh Shetye [^1] footnotemark: 1 Harshit Gupta2Manav Chaudhary3Manish Shrivastava1Vasudeva Varma1 1IIIT Hyderabad  2Microsoft, India  3Sentisum {maitreya.chitale, ketaki.shetye}@research.iiit.ac.in guptaharshi@microsoft.com manav@sentisum.com {m.shrivastava, vv}@iiit.ac.in
1 Introduction
Recent advancements have focused on assisting reviewers and directly generating peer reviews Yu et al. (2024); Aitymbetov and Zorbas (2025); D’Arcy et al. (2024); Idahl and Ahmadi (2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) in this domain Ye et al. (2024). Relying on LLMs to generate or inform final review decisions remains risky due to issues of bias, overconfidence, and superficial feedback and these decision-oriented tools lack utility for authors in the pre-submission phase. Consequently, we advocate for a new paradigm towards author-centric support. We introduce AutoRev, an author-centric system that generates structured, actionable pre-submission feedback containing Summary, Strengths, Weaknesses, and Questions, without producing acceptance recommendations or replacing human peer review.
@@ 1 Introduction
Figure 1: Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.
@@
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community Chitale et al. (2025); Gao et al. (2025); Agarwal et al. (2025); Zhu et al. (2025). Inspired by this, we leverage the graph’s capability as an extractor of passages to significantly reduce the input token length to the LLM. Reducing the token length of the input is shown to enhance LLM performance Hsieh et al. (2024). To summarize our main contributions:
  1. 1.
    We propose an author-centric system for generating concise, structured pre-submission feedback for academic papers.
  2. 2.
    We introduce a graph-based retrieval framework that learns to identify salient passages, reducing LLM input length while improving feedback quality.
  3. 3.
    We validate our approach using comprehensive human evaluations to assess feedback quality and factual grounding, in addition to standard automatic metrics.
2 Leveraging Peer Reviews as Proxy for Pre-Submission Feedback
In the absence of large-scale, publicly available datasets explicitly designed for pre-submission feedback, we use human-written peer reviews as a proxy, as they provide expert critiques of writing, methodology, and presentation aligned with the guidance authors seek prior to submission. In particular, the Weaknesses and Questions sections offer actionable critiques and clarifications. Since the dataset spans papers at varying stages of writing maturity, the associated reviews reflect feedback targeted at authors with differing levels of experience. We further support this choice through an impact-oriented user study (Section 6.3.4).
3 Related Works
Recent work has explored the use of NLP techniques to support scholarly writing and review workflows, including reviewer selection Aitymbetov and Zorbas (2025), meta-review generation Kumar et al. (2023), and automated review generation Yu et al. (2024). Many of these approaches rely on large language models, ranging from modular agent-based systems D’Arcy et al. (2024) to fine-tuned review generators Idahl and Ahmadi (2025). In parallel, a separate line of work focuses on author-assistive tools for specific writing sub-tasks, such as limitation suggestion Faizullah et al. (2024), citation recommendation and literature discovery Gu and Hahnloser (2023), and drafting individual sections like related work Hu and Wan (2014) or introductions Liebling et al. (2025). While these are effective within their scoped objectives, they are either reviewer-facing or narrowly targeted, and often operate on full-paper inputs at inference time. As a result, they raise concerns around hallucination, superficial feedback, computational cost, and overconfidence in model-generated judgments Ye et al. (2024), while failing to provide holistic, structured feedback that supports authors prior to submission.
Yu et al. (2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4 OpenAI et al. (2024) consolidation and proposes a self-correction strategy during fine-tuning. We adopt SEA as our baseline as it is trained on the publicly available ICLR 2024 dataset and generates the same structured fields as AutoRev, among others.
@@ 3 Related Works
Studies have explored graph-based extensions to Retrieval-Augmented Generation (RAG) Lewis et al. (2020) to improve grounding and reasoning in LLMs. Zhu et al. (2025) provide a comprehensive taxonomy of Graph-RAG systems based on database construction, retrieval mechanisms, and processing pipelines. Within this framework, AutoRev represents a domain-specific Graph-RAG approach that constructs a hierarchical graph directly from a scientific paper and employs a learning-based Graph Attention Network veličković2018graph for passage retrieval.
4 Dataset
@@
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4-consolidated reviews for NeurIPS 2023 are not available. Hence, our analysis focuses only on the ICLR 2024 papers. Table 1 summarizes statistics for the ICLR 2024 papers Yu et al. (2024).
@@
Table 1: Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
Each paper is associated with multiple reviews containing sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating. We retain only the Summary, Strengths, Weaknesses, and Questions sections, as these encode constructive feedback for authors-centric pre-submission support (see Section 2), and exclude rating-based fields that correspond to final decision-making. For training, we use the GPT-4 consolidated reviews released by Yu et al. (2024), which merge multiple human-written reviews into a single structured review per paper.
To assess cross-venue performance, we curate a supplementary dataset of 10 papers from COLM and NeurIPS 2025 (CNT_10). This set complements the ICLR 2024 test data (ICT) for the human evaluation detailed in Section 6.3.
Figure 2: The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to structured feedback generation. The trained GNN selects important passages without requiring human-generated feedback. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final structured feedback, without external grounding.
5 Methodology
In this section, we describe the implementation of our proposed framework. Section 5.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) (Karpukhin et al., 2020). The extracted passages are used as training data in Section 5.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated feedback, as discussed in Section 5.4. Post that, we describe our process of fine-tuning LLMs to generate the final structured feedback, in Section 5.5. An overview of our methodology can be found in Figure 2.
5.1 DPR-based Passage Extraction
Dense Passage Retrieval (DPR) is a neural retrieval model trained primarily on question-answering datasets that encodes queries and passages into dense vector representations. Given an input query, DPR retrieves relevant passages based on semantic similarity in the embedding space. We use a pre-trained DPR model to identify passages corresponding to the Summary, Strengths, Weaknesses, and Questions sections of reviews for each paper in the training set. Specifically, we use the context encoder [^1] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and query encoder [^2] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.